How many kj of energy are required to force 115g of ice at -15c to turn into steam at 105c? The molar mass of 18g/mol
1
Expert's answer
2020-04-13T02:33:05-0400

molar heat of fusion, "Q=N*C*\delta T"

"N=m\/RMM"

N=115/18=6.388moles

Q=6.388*4200*(105--15)=3219552J

Q=3219.55kJ